OpenStudy (anonymous):

The center of the circle with equation (x - 3)2 + (y + 2)2 = 4 has coordinates ___.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow (x - h)^2 + (y - b)^2 = r^2 }\) (h,b) are the center co-ordinates.

OpenStudy (anonymous):

can you please tell me the answer lol i have no idea what you just wrote

OpenStudy (anonymous):

look at corresponding points

Parth (parthkohli):

Hmm 3 = h and b = -2 (3,-2) are the center co-ordinates

Parth (parthkohli):

Please do others yourself

OpenStudy (anonymous):

thank you!